OpenStudy (anonymous):

Suppose that f is a function with f2=-1 and f'2=-32. If hx=1/f(x)^12 then h'2=

OpenStudy (ash2326):

We have \[ h(x)=\frac{1}{{f(x)}^{12}}\] Let's find h'(x) \[ h'(x)= -12 \times {f(x)}^{-13} \times f'(x)\] h'(2), let's substitute x=2 in h'(x) \[ h'(2)= -12 \times {f(2)}^{-13} \times f'(2)\] We have f(2)=-1 and f'(2)=-32 \[ h'(2)= -12 \times {-1}^{-13} \times -32\] we know \({(-1)}^{-13}=-1\) so \[h'(2)=-12\times -1 \times -32= -384\]

OpenStudy (anonymous):

thanks i got it now

OpenStudy (ash2326):

Welcome dd